slope=-1/5, passing through (1,2)
y-(-2)=-1/5(x-1)
point slope form: y+2=-1/5(x-1)
translate into slope-intercept form: This is where I get lost.

Answer by Alan3354(69443)

slope=-1/5, passing through (1,2)
y-(-2)=-1/5(x-1) ***** y-2
point slope form: y-2 = -1/5(x-1)
translate into slope-intercept form: This is where I get lost.
Just solve for y
y-2 = -1/5(x-1)
y = -1/5(x-1) + 2
y = -x/5 + 1/5 + 2
y = -x/5 + 11/5